OpenStudy (anonymous):

A ball of clay has a mass of 2.65 pounds and a volume of 0.5025 quarts. What is the density in grams per milliliter?

OpenStudy (anonymous):

1 pound = 453.59 g 1 US quart = 946.35 ml so: 2,65 * 453,59 = ? g and 0,5025 * 946,35 = ? ml and density is: rho= m/V so: rho = ? g/ ? ml = ? g/ml